# Near-field Beam Training with Sparse DFT Codebook

### v1

@@ V-C Phase 3: Beam Sweeping with Polar-domain Codebook
6:  Activate a central sub-array with $M$ antennas and use codebook ${{\mathcal{V}}}_{\rm{Sub}}$ to examine $U$ candidate angles estimated in the first phase.
@@ V-D Discussions
Beam training overhead: In the first phase, the overhead is $Q=\frac{N-1}{U}+1$. Moreover, the overhead of the second and third phases are $U$ and $V$, respectively. Finally, the overall overhead of the proposed beam training scheme is $T^{\rm(3P)}=Q+U+V$. Given $Q=\frac{N-1}{U}+1\in\mathbb{N}$ with $\mathbb{N}$ denoting the positive integer set, the beam training overhead of the proposed method enabled by the sparse DFT codebook can be recast as
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ U\in\mathcal{F}\triangleq\{\frac{N-1}{U}+1\in\mathbb{N},\leavevmode\nobreak\ U\in\mathbb{N}\}.$ (37b)
@@ VI-A System Setup and Benchmark Schemes
Figure 9: Effective rate versus SNR.
@@
Figure 10: Achievable rate versus user range.
@@ VI-B Performance Analysis
Figure 13: Achievable rate versus number of antennas.
@@ VII Conclusion
In this paper, we proposed a novel near-field beam training scheme enabled by the sparse DFT codebook (sparse far-field beamforming vectors) to construct periodic received beam pattern at the user. To this end, we have shown that the angular periodicity of the received beam pattern boosts reduction in sweeping space, thereby significantly decreasing beam training overhead. Specifically, the middle of the angular support within a period contains the whole user angle information. Then, an activated central sub-array can resolve the angular ambiguity followed by polar-domain codebook sweeping in the best user angle. Finally, numerical results were presented to show that the proposed beam training scheme can achieve nearly the same performance in the high-SNR regime with extremely low overhead.

### v2

@@ V-C Phase 3: Beam Sweeping with Polar-domain Codebook
The detailed procedures of the proposed three-phase beam training method are summarized in Algorithm 1.
@@
6:  Activate a central subarray with $M$ antennas and use codebook ${{\mathcal{V}}}_{\rm{Sub}}$ to examine $U$ candidate angles estimated in the first phase.
@@ V-D Discussions
Beam training overhead: In the first phase, the overhead is $Q=\frac{N-1}{U}+1$. Moreover, the overhead of the second and third phases are $U$ and $V$, respectively. Finally, the overall overhead of the proposed beam training scheme is $T^{\rm(3P)}=Q+U+V$. Given $Q=\frac{N-1}{U}+1\in\mathbb{Z}^{+}$ with $\mathbb{Z}^{+}$ denoting the positive integer set, the beam training overhead of the proposed method enabled by the sparse DFT codebook can be recast as
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ U\in\mathcal{F}\triangleq\{\frac{N-1}{U}\!+\!1\in\mathbb{Z}^{+},U\in\mathbb{Z}^{+}\}.$ (39b)
@@ VI-A System Setup and Benchmark Schemes
Figure 9: Achievable rate versus user range.
@@ VI-B Performance Analysis
Figure 13: Achievable rate versus the number of antennas in the central subarray.
@@
Figure 14: Achievable rate versus number of antennas.
In Fig. 13, we plot the achievable rate versus the number of antennas in the central subarray. It can be observed that the achievable rate of the proposed scheme suffers from significant performance loss when the number of antennas in the central subarray is sufficiently small or large. This can be explained by two facts: 1) When $M$ is sufficiently large (Violation of Criteria 1 in Section V-B), the user are more likely to be located in the near-field region of the central subarray, for which the energy-spread effect is dominant and thus results in significant performance loss; 2) When $M$ is sufficiently small (Violation of Criteria 2 in Section V-B), the beam width of the central subarray becomes too large, introducing extra ambiguity and thus failing to distinguish candidate user angles.
@@ VII Conclusion
In this paper, we proposed a novel near-field beam training scheme enabled by the sparse DFT codebook (sparse far-field beamforming vectors) to construct periodic received beam pattern at the user. To this end, we showed that the angular periodicity of the received beam pattern boosts reduction in sweeping space, thereby significantly decreasing beam training overhead. Specifically, the middle of the angular support within a period contains the user angle information. Then, an activated central subarray can resolve the angular ambiguity followed by polar-domain codebook sweeping in the best user angle. Finally, numerical results were presented to show that the proposed beam training scheme can achieve nearly the same performance in the high-SNR regime with the exhaustive-search scheme, while significantly reducing the beam training overhead.